5. A motorist drives 100 km from one city to another in 2.60 h but makes the return trip in only 1.90 h. What is the average speed for the total trip?
Answer #1

Average speed is the speed with which the motorist travelled the entire journey.

Average speed = Total distance / Total time

Total distance = 100 + 100 = 200 km

Total time = 2.6 + 1.9 = 4.5 hr

Average speed = 200/4.5 = 44.44 km/hr